please give an explanation!
(4 pts) A heterozygous genetic condition called creeper in chickens produces shortened and deformed legs and wings, giving
Answer #1

Given, 2 775 Normal = 388 ... Total = 775+388=1163. The Ratio b/w creeper normal = 3:1 Chi square Expecteel categary I Obern